OpenStudy (anonymous):

integrate dx/1+2sin^x from 0 to pie

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@genny7

OpenStudy (anonymous):

Do you mean\[\int\frac{dx}{1+2\sin^2x}~~?\]

OpenStudy (anonymous):

its definite integration from 0 to pie

OpenStudy (anonymous):

Yes I get that (but omitted it), but the integrand you provided doesn't make sense.

OpenStudy (anonymous):

i didnt understand

OpenStudy (anonymous):

What do you mean by \(\sin^x\)? What is the argument?

OpenStudy (anonymous):

its sin^2x

OpenStudy (anonymous):

So it is \[\int_0^\pi\frac{dx}{1+2\sin^2x}~~?\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

divide top and bottom by cos^2x

ganeshie8 (ganeshie8):

\[\large \large \int \dfrac{dx}{1+2\sin^2x} = \int \dfrac{\sec^2x~dx}{1+3\tan^2x}\]

ganeshie8 (ganeshie8):

next substitution is quite obvious..

OpenStudy (anonymous):

Another way to do it is dividing by \(\sin^2x\) in place of \(\cos^2x\). Same process in the end.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how limit will change

ganeshie8 (ganeshie8):

Ahh looks thats going to be tricky we need to split the integral i guess

OpenStudy (anonymous):

how tell me

OpenStudy (anonymous):

i m not getting

ganeshie8 (ganeshie8):

@sidsiddhartha

ganeshie8 (ganeshie8):

@aum

ganeshie8 (ganeshie8):

dividing by sin^2x and substituting cotx works out without complications as cotx is continuous in (0,pi) but dividing by cos^2x and tanx substitution also must work if we split the integral, right ?

OpenStudy (anonymous):

is it an odd function

ganeshie8 (ganeshie8):

not sure if i am making sense, im bit confused hmm

OpenStudy (aum):

This is an even function. So the integral could be 2 * [0, pi/2].

OpenStudy (anonymous):

ok

OpenStudy (aum):

oops sorry. I read the limits wrong.

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

going for lunch brb, thanks aum :)

OpenStudy (anonymous):

someone please help me

OpenStudy (anonymous):

\[\begin{align*}\int\frac{dx}{1+2\sin^2x}&=\int\frac{\csc^2x}{\csc^2x+2}~dx\\\\ &=\int\frac{\csc^2x}{(1+\cot^2x)+2}~dx\\\\ &=\int\frac{\csc^2x}{\cot^2x+3}~dx\\\\ &=-\int\frac{du}{u^2+3}&\text{setting }u=\cot x\\&&\text{so }-du=\csc^2x~dx\\\\ &=-\int\frac{\sqrt3\sec^2t}{(\sqrt3\tan t)^2+3}~dt&\text{setting }u=\sqrt 3\tan t\\ &&\text{so }du=\sqrt 3\sec^2t~dt\\\\ &=-\int\frac{\sqrt3\sec^2t}{3\tan^2 t+3}~dt\\\\ &=-\frac{\sqrt{3}}{3}\int\frac{\sec^2t}{\tan^2 t+1}~dt\\\\ &=-\frac{\sqrt{3}}{3}\int dt \end{align*}\] Integrate and back-substitute.

OpenStudy (anonymous):

i got the answer.by another method just now

ganeshie8 (ganeshie8):

looks you need to use some symmetry http://math.stackexchange.com/questions/958920/evaluate-int-limits-0-pi-dfracdx12-sin2x whats your other method ?

OpenStudy (anonymous):

thanks ganeshie bt i got the answer

ganeshie8 (ganeshie8):

yeah i am stuck at understanding the bounds because F(pi) - F(0) gives you 0. may i knw your other method ?

OpenStudy (anonymous):

i break the integation in two parts from o to pie/2 and pie/2 to pie

OpenStudy (anonymous):

i did this when i got sec^2dx/1+3tan^2x

ganeshie8 (ganeshie8):

you will get 0 right ?

ganeshie8 (ganeshie8):

F(pi/2) - F(0) + F(pi) - F(pi/2) F(pi) - F(0) 0

OpenStudy (anonymous):

no pie/root 3

ganeshie8 (ganeshie8):

splitting is giving me 0. see above

OpenStudy (anonymous):

limit will change from0 to infinity and -infnity to 0

ganeshie8 (ganeshie8):

you're right ! splitting will do (: